OpenStudy (anonymous):

Solve the equation for y 4x -5y = 9 A. Y = 5/4x + 9/4 B. Y = -4x - 9 C. Y = 4/5x - 9/5 D. Y = -5/4x - 5/9

OpenStudy (anonymous):

Make the equation \(4x-5y=9\) to \(y=mx+b\).

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

\(\huge 4x-5y=9\) Subtract \(4x\) on both sides. \( -4x+4x-5y=-4x+9\) Then you'll get \(\huge -5y=-4x+9\) Now, divide \(-5\) on both sides. \(\huge \frac{-5y}{-5}=\frac{-4x}{5}+\frac{9}{5}\).

OpenStudy (anonymous):

Okay thank you

OpenStudy (anonymous):

So, \(\huge y=\frac{-4x}{-5}+\frac{9}{-5}\) OR .\(\huge y=\frac{4x}{5}-\frac{9}{5}\) :D

OpenStudy (anonymous):

You're welcome! Just a correction: it should be -5 in my second reply. :)